OpenStudy (anonymous):

All 28 members in Crestview High School’s Ski Club went on a one-day ski trip. Members can rent skis for $16.00 per day or snowboards for $19.00 per day. The club paid a total of $478 for rental equipment. How many members rented skis and how many rented snowboards? a. 10 members rented skis 18 members rented snowboards c. 18 members rented skis 10 members rented snowboards b. 19 members rented skis 9 members rented snowboards d. 20 members rented skis 8 members rented snowboards @skinny23

OpenStudy (anonymous):

D

OpenStudy (anonymous):

18 members rented ski 10 members rented snowboards 18*16=288 10*19=190 288+190= 478